# backgroundcolor=, commentstyle=, breakatwhitespace=false, keywordstyle=, language=Java, stringstyle=, basicstyle=, showstringspaces=false

### v1

@@ 6.1 Implications
  • The results in RQ3 show that the PLMs we considered achieve mediocre results when detecting TD in source code. This implies that there is still room for improvement: more advanced detection models are needed. We anticipate that LLMs could be an eventual solution to this problem, as they have been trained with a huge amount of data including source code, with the potential to better capture the intrinsic features of TD contained in source code. This, however, needs further refinement and empirical evidence and is part of our future work.
@@ 7 Related Work
Technical debt detection datasets: The availability of datasets is critical for advancing research in technical debt related tasks, providing empirical evidence to validate the effectiveness of detection techniques. Large-scale studies of TD frequently focus on a specific category, with Self-Admitted Technical Debt (SATD) being a prevalent area of focus. Maldonado et al. [42, 16] presented a widely utilized dataset, comprising more than 62,000 comments from 10 Java projects, classified into five types of TD and non-SATD. Instead of using only comment, Li et al. [37] introduced the dataset by investigating four different sources: code comments, commit messages, pull requests, and issue tracking systems. In addition, they merged code debt and design debt into a single category due to their high similarity. To explore programming languages beyond Java, Sharma et al. [58] introduced a dataset for examining SATD in the R language, which includes over 140,000 samples and expands the number of TD types to 12. Furthermore, several datasets [47, 41, 45, 30] have been introduced to address code smells, which is a related issue to TD.
@@
SATD detection techniques: Early methods for identifying and detecting SATD relied on rule-based approaches that searched for matching keywords or phrases [26, 62]. For example, the Matches Task Annotation Tags (MAT) [26] method demonstrated that simply matching a set of commonly used task annotation tags, such as TODO, FIXME, HACK, and XXX, can achieve a significant performance in identifying SATD. In PENTACET, Sridharan et al. [62] built upon the 64 SATD detection patterns initially introduced by Potdar and Shihab [48]. By leveraging the Sense2Vec tool, they expanded the set to 1,041 patterns, significantly increasing the scope for identifying SATD. More advanced methods have been proposed using machine learning approaches [42, 16, 55], where classifiers like maximum entropy and Naive Bayes multinomial classifiers are trained to detect various types of TD. Recently, supervised deep learning methods have been introduced, demonstrating superior performance compared to rule-based and traditional machine learning approaches. Li et al. [37] introduced a method called MT-Text-CNN, which utilizes a convolutional neural network combined with multi-task learning to detect SATD across multiple sources. Meanwhile, Yu et al. [77] proposed a method utilizing bidirectional long short-term memory (BiLSTM) networks with an attention mechanism and a balanced cross-entropy loss function to mitigate the imbalance problem in SATD identification. Besides, several approaches have utilized pretrained language models, achieving state-of-the-art performance in SATD-related tasks. Sharma et al. [58] demonstrated that ALBERT [31] and RoBERTa [39] significantly outperform traditional machine learning and CNN-based methods in identifying and classifying SATD categories in the R language. Furthermore, Sheikhari et al. [59] illustrated the superior performance of the LLM model Flan-T5 [12] across three SATD-related tasks.
8 Conclusions and Future Work
@@
Acknowledgments
This work has been partially supported by the EMELIOT national research project, which has been funded by the MUR under the PRIN 2020 program (Contract 2020W3A5FY). The work has been also partially supported by the European Union–NextGenerationEU through the Italian Ministry of University and Research, Projects PRIN 2022 PNRR “FRINGE: context-aware FaiRness engineerING in complex software systEms” grant n. P2022553SL. We acknowledge the Italian “PRIN 2022” project TRex-SE: “Trustworthy Recommenders for Software Engineers,” grant n. 2022LKJWHC. Our research is also funded by Hanoi University of Science and Technology (HUST), Vietnam under project number T2023-PC-002.

### v2

@@ 6.1. Implications
  • The results in RQ2 show that the PLMs we considered achieve mediocre results when detecting TD in source code. This implies that there is still room for improvement: more advanced detection models are needed. We anticipate that LLMs could be an eventual solution to this problem, as they have been trained with a huge amount of data including source code, with the potential to better capture the intrinsic features of TD contained in source code. This, however, needs further refinement and empirical evidence and is part of our future work.
@@ 7. Related Work
Technical debt detection datasets: The availability of datasets is critical for advancing research in technical debt related tasks, providing empirical evidence to validate the effectiveness of detection techniques. Large-scale studies of TD frequently focus on a specific category, with Self-Admitted Technical Debt (SATD) being a prevalent area of focus. Maldonado et al. (Maldonado and Shihab, 2015; da Silva Maldonado et al., 2017) presented a widely utilized dataset, comprising more than 62,000 comments from 10 Java projects, classified into five types of TD and non-SATD. Instead of using only comment, Li et al. (Li et al., 2023) introduced the dataset by investigating four different sources: code comments, commit messages, pull requests, and issue tracking systems. In addition, they merged code debt and design debt into a single category due to their high similarity. To explore programming languages beyond Java, Sharma et al. (Sharma et al., 2022) introduced a dataset for examining SATD in the R language, which includes over 140,000 samples and expands the number of TD types to 12. Furthermore, several datasets (Palomba et al., 2015; Madeyski and Lewowski, 2020; Nandani et al., 2023; Kovačević et al., 2024) have been introduced to address code smells, which is a related issue to TD.
@@
SATD detection techniques: Early methods for identifying and detecting SATD relied on rule-based approaches that searched for matching keywords or phrases (Guo et al., 2021; Sridharan et al., 2023). For example, the Matches Task Annotation Tags (MAT) (Guo et al., 2021) method demonstrated that simply matching a set of commonly used task annotation tags, such as TODO, FIXME, HACK, and XXX, can achieve a significant performance in identifying SATD. In PENTACET, Sridharan et al. (Sridharan et al., 2023) built upon the 64 SATD detection patterns initially introduced by Potdar and Shihab (Potdar and Shihab, 2014). By leveraging the Sense2Vec tool, they expanded the set to 1,041 patterns, significantly increasing the scope for identifying SATD. More advanced methods have been proposed using machine learning approaches (Maldonado and Shihab, 2015; da Silva Maldonado et al., 2017; Sala et al., 2021), where classifiers like maximum entropy and Naive Bayes multinomial classifiers are trained to detect various types of TD. Recently, supervised deep learning methods have been introduced, demonstrating superior performance compared to rule-based and traditional machine learning approaches. Li et al. (Li et al., 2023) introduced a method called MT-Text-CNN, which utilizes a convolutional neural network combined with multi-task learning to detect SATD across multiple sources. Meanwhile, Yu et al. (Yu et al., 2021) proposed a method utilizing bidirectional long short-term memory (BiLSTM) networks with an attention mechanism and a balanced cross-entropy loss function to mitigate the imbalance problem in SATD identification. Besides, several approaches have utilized pretrained language models, achieving state-of-the-art performance in SATD-related tasks. Sharma et al. (Sharma et al., 2022) demonstrated that ALBERT (Lan et al., 2019) and RoBERTa (Liu et al., 2019) significantly outperform traditional machine learning and CNN-based methods in identifying and classifying SATD categories in the R language. Furthermore, Sheikhari et al. (Sheikhaei et al., 2024) illustrated the superior performance of the LLM model Flan-T5 (Chung et al., 2024) across three SATD-related tasks. Some methods incorporate corresponding source code to support SATD detection (Russo et al., 2022; Zampetti et al., 2020; Russo et al., 2025). For example, Russo et al. (Russo et al., 2022) introduced WeakSATD to analyze C source code from Chromium projects, aiming to determine whether code blocks linked to SATD comments contain potential weaknesses. Meanwhile, VulSATD (Russo et al., 2025) leverages CodeBERT to jointly encode comment–code pairs in a multi-task setup, with separate output heads for identifying SATD and detecting code vulnerabilities.
8. Conclusions and Future Work
@@
Acknowledgements.
This paper has been partially supported by the MOSAICO project (Management, Orchestration and Supervision of AI-agent COmmunities for reliable AI in software engineering) that has received funding from the European Union under the Horizon Research and Innovation Action (Grant Agreement No. 101189664). The work has been partially supported by the EMELIOT national research project, which has been funded by the MUR under the PRIN 2020 program (Contract 2020W3A5FY). It has been also partially supported by the European Union–NextGenerationEU through the Italian Ministry of University and Research, Projects PRIN 2022 PNRR “FRINGE: context-aware FaiRness engineerING in complex software systEms” grant n. P2022553SL. We acknowledge the Italian “PRIN 2022” project TRex-SE: “Trustworthy Recommenders for Software Engineers,” grant n. 2022LKJWHC. Our research is also funded by Hanoi University of Science and Technology (HUST), Vietnam under project number T2023-PC-002.